OpenStudy (anonymous):

help....CHEMISTRY.... Work the entire problem from the beginning. How many grams of H2 would be formed if 34 grams of carbon reacted with an unlimited amount of H2O? The reaction is: C + H2O → CO + H2 The atomic mass of C is 12.01 g/mole. The atomic mass of H2 is 2.016 g/mole. grams of hydrogen 823 5.7 14

OpenStudy (aaronq):

It's a simple stoichometry problem: Start by converting the mass of carbon to moles using: \(n=\dfrac{m}{M}\) The follow this: -------------------------------------------------- outline: Using the stoichiometric coefficients to find moles produced. e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can solve for moles of product. -------------------------------------------------------- For example: if you have moles of B= 2, how many moles of C did you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_c=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To convert moles back to mass, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.

OpenStudy (preetha):

Aaron, this is a masterly explanation. I would like to frame this and display it! Thanks

OpenStudy (aaronq):

Thank you. it's something i wrote for someone else before so i just "copy and paste it".

OpenStudy (anonymous):

So the answer is?